### v5

@@ 1 Introduction
We instantiate this principle in the context of stochastic weakly convex objectives, a broad class of nonsmooth, nonconvex functions encompassing many widely studied models [3, 4, 5]. Our main application is to stochastic convex-composite objectives [6, 7], which include robust phase retrieval [8], blind deconvolution [9], matrix sensing [10, 11], conditional Value-at-Risk [12], and nonlinear quantile regressions [13]. These models appear across imaging, statistics, machine learning, and risk management. For this class, we derive uniform convergence rates for subdifferentials by studying the convergence behavior of stochastic subgradient mappings (Theorem 4 and Theorem 5). Our analysis complements prior work by (i) removing differentiability assumptions on the population risk, and (ii) establishing tight convergence rates under the Hausdorff distance, a topologically weaker notion than graphical convergence that more directly captures the geometry of subdifferential sets. For instance, in the setting of robust phase retrieval [8, 14], we establish a sharp convergence rate of $\sqrt{d/m}$ (modulo logarithmic factors on $d,m$) for the uniform convergence of subdifferential mappings under the Hausdorff distance, where $d$ is the problem’s dimension and $m$ is the sample size (Corollary 1). This is complementary to the state-of-the-art $4]{d/m}$ rate, which is established under the graphical distance of subdifferentials using an alternative approach based on Attouch’s epigraphical convergence theorem [15, 16]. Our results lead to new understanding into nonsmooth empirical risk landscapes, including how the locations of stationary points evolve with sample size—providing a sharper understanding of finite-sample behavior in high-dimensional stochastic optimization.
@@ 2 Notation and Basic Definitions
[…]
@@ 2.2 Set-valued Analysis
[…]
@@ 2.4 Measurability Issues
Notably, $\mathbb{P}^{*}$ is subadditive: $\mathbb{P}^{*}(B_{1}\cup B_{2})\leq\mathbb{P}^{*}(B_{1})+\mathbb{P}^{*}(B_{2})$, following the union bound. Correspondingly, the inner measure is $\mathbb{P}_{*}(B)=\sup\{\mathbb{P}(A)\colon B\supseteq A,A\in\mathscr{G}\}$. Clearly, $\mathbb{P}_{*}(B)=1-\mathbb{P}^{*}(\Xi\backslash B)$. Finally, when $B\in\mathscr{G}$ is measurable, then the inner and outer measure agree: $\mathbb{P}^{*}(B)=\mathbb{P}_{*}(B)=\mathbb{P}(B)$.
@@ 3.1 Uniform Subdifferential Bounds via Selections
Here, $D_{i}$ denotes the set of differentiable points of $f_{i}$, which is dense in the open set $\mathcal{O}$ (meaning that $\mathcal{O}\backslash D_{i}$ has Lebesgue measure 00), and ${\rm conv}$ represents the convex hull of a set. By definition, the sets $\mathcal{G}_{i}(x)$ are closed at every $x$, and the set-valued mappings $x\mapsto\mathcal{G}_{i}(x)$ are locally bounded as the functions $f_{i}$ are locally Lipschitz.
@@
We now argue that we can, without loss of generality, assume that for every $k$, $x_{k}\in D_{2}$ (the set of differentiable points of $f_{2}$) as well. If some $x_{k}\not\in D_{2}$, applying Clarke’s subdifferential characterization to $f_{1}$ at $x_{k}$ (3.2), we can replace such $x_{k}$ with some $x_{k}^{\prime}$ belonging to the subset $D_{1}\cap D_{2}$ dense in $\mathcal{O}$ (note $\mathcal{O}\backslash(D_{1}\cap D_{2})$ has Lebesgue measure 00) such that $\left\|{x_{k}^{\prime}-x_{k}}\right\|\to 0$ and $\left\|{\nabla f_{1}(x_{k}^{\prime})-\nabla f_{1}(x_{k})}\right\|\to 0$ as $k\to\infty$. As a result, $x_{k}^{\prime}$ preserves all the properties of $x_{k}$, including $x_{k}^{\prime}\to x$, $x_{k}^{\prime}\in D_{1}$ and $\nabla f_{1}(x_{k}^{\prime})\to v_{1}^{*}$, while guaranteeing that $x_{k}^{\prime}\in D_{2}$ for all $k$.
@@
While not essential to the rest of the paper, our proof of Theorem 1 naturally extends to the following generalization, Theorem 1, which could be of independent interest to specialists in variational analysis. To state it precisely, we recall the definition of the Clarke subdifferential of a function $f$ from the standard reference [17, equation 8(32)].
Theorem 1.
Let $f_{1}$ and $f_{2}$ be locally Lipschitz functions from $\mathcal{O}$ to $\mathbb{R}$, where $\mathcal{O}$ is an open set in $\mathbb{R}^{d}$. Let ¯$\bar{\partial}f_{1}$ and ¯$\bar{\partial}f_{2}$ denote the Clarke subdifferential of $f_{1}$ and $f_{2}$, respectively. For every pair of selections $g_{1}$ and $g_{2}$ of the subdifferentials ¯$\bar{\partial}f_{1}$ and ¯$\bar{\partial}f_{2}$ on $\mathcal{O}$, respectively, i.e., obeying $g_{1}(x)\in\bar{\partial}f_{1}(x)$ and $g_{2}(x)\in\bar{\partial}f_{2}(x)$ for all $x\in\mathcal{O}$, the following inequality holds:
$\sup_{x\in\mathcal{O}}\mathbb{H}(\bar{\partial}f_{1}(x),\bar{\partial}f_{2}(x))\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$ (3.10)
Proof.
By [17, Theorem 8.49 + Theorem 9.13], ¯$\bar{\partial}f_{i}(x)$ is a convex and compact set at every $x\in\mathcal{O}$. In fact, by Rademacher’s theorem [17, Theorem 9.61], ¯$\bar{\partial}f_{i}(x)$ obeys the characterization:
¯$\bar{\partial}f_{i}(x)={\rm conv}\left\{\lim_{k\to\infty}\nabla f_{i}(x_{k})\colon x_{k}\in D_{i},x_{k}\to x,\lim_{k\to\infty}\nabla f_{i}(x_{k})~{}\text{exists}\right\}.$
Given these key ingredients, the proof of Theorem 1 follows verbatim from that of Theorem 1, with the sole exception that $\partial f_{i}$ is replaced by ¯$\bar{\partial}f_{i}$. ∎
@@ 3.3 A Class of Stochastic Weakly Convex Minimization Problems
We now revisit the conditions on the convex function $h$ and smooth function $c(x,\xi)$ that guarantees the weak-convexity assumptions for the composition $f(x,\xi)=h(c(x,\xi))$, ensuring the regularity required for subdifferential calculus rules. We take the following conditions—local Lipschitzian and integrability conditions on $h,c$—from [7, Section 2.2].
@@ 4.1 Subgradient Selections
where $z\mapsto\mathbf{1}(z\geq t_{j})$ is an indicator function that is a subgradient for the mapping $z\mapsto(z-t_{j})_{+}$, taking the value $1$ when $z\geq t_{j}$ and 00 otherwise. Combining these, we define a subgradient $g$ for $h$ at every $z$:
@@ 4.2 A Uniform Bound on Subgradient Selections
=$\mathcal{F}=\{\{\xi\in\Xi\colon c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}.$ (4.12)
@@ 4.3 Implication: A Uniform Bound on Subdifferentials
In the literature on uniform convergence rates of subdifferential mappings for stochastic weakly-convex objectives, methods using Moreau envelope smoothing [26] and Attouch’s epigraphical convergence theorem [15] achieve a convergence rate of $4]{1/m}$ suboptimal in terms of the sample size $m$ [14, 16]. This $4]{1/m}$ convergence is measured using graphical distance, a notion which is topologically stronger than the uniform convergence criterion we consider in Theorem 5 (see Remark of Theorem 1). Our work complements these studies by showing a convergence rate’s dependence on sample size $m$ to be $\sqrt{1/m}$ (modulo logarithmic factors) under a topologically weaker notion of convergence.
@@ 5 VC Dimension Bounds
=$\mathcal{F}=\{\{\xi\in\Xi\colon c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}$
@@
=$\begin{split}\mathcal{B}&=\{(b_{1},b_{2},\ldots,b_{N})\in\{0,1\}^{N}\colon\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{there exists $z\in\mathbb{R}^{d+1}$ such that}~{}b_{i}=\mathbf{1}\{q_{i}(z)\geq 0\}~{}\text{for all $1\leq i\leq N$}\}.\end{split}$
@@ 6.1 Illustration I: Phase Retrieval
as 00 is the only nondifferentiable point of $h$. The corresponding $\mathcal{F}$ is given by
=$\mathcal{F}=\left\{\{a\in\mathbb{R}^{d},b\in\mathbb{R}\colon\langle a,x\rangle^{2}-b\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\right\}.$
@@ 6.1.1 Noiseless Phase Retrieval
For standard normal measurement vectors $a\sim\mathsf{N}(0,I)$, a recent important and interesting study first characterizes $\mathcal{Z}$, the locations of stationary points of the population objective $\Phi(x)$ [14]. Remarkably, the authors further develop a quantitative version of Attouch’s epi-convergence theorem to show that the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt[4]{d/m}$14]. More precisely, [14, Theorem 5.2, Corollary 6.3] prove that, there exist numerical constants $c,C>0$ such that when $m\geq Cd$, with an inner probability at least $1-C\exp(-cd)$ (recall the definition of inner probability from Section 2.4):
@@ 7 Discussion
  • Can Theorem 1 be utilized to deduce concrete convergence rates for other interesting class of stochastic locally Lipschitz minimization problems? A perhaps interesting example is the convergence of subdifferentials in stochastic difference-of-convex optimization problems [58].
@@
Acknowledgements Feng Ruan extends his gratitude to his friends and advisor—John Duchi, Qiyang Han, X.Y. Han, Kate Yao and Yichen Zhang—for their support and patience as he tackled this once seemingly challenging problem. A shift in perspective during his time at Northwestern led him to this solution. He is also grateful to Damek Davis and Dmitriy Drusvyatskiy for their valuable feedback on the initial manuscript draft posted on arXiv. Feng Ruan would also like to express gratitude to Alexander Shapiro, and the anonymous reviewers for their valuable discussions on literature and insights regarding the proof of Theorem 1 using variational analytic tools.
@@ Appendix A Proof of Theorem 4
$\mathbb{P}^{*}\left(\sup_{z_{1}\in\mathcal{Z}_{1},z_{2}\in\mathcal{Z}_{2},t\in\mathbb{R}}\left\|{K_{m}(z_{1},z_{2},t)-K(z_{1},z_{2},t)}\right\|\geq C\sigma r\cdot\max\{\Delta_{\mathcal{W}},\Delta_{\mathcal{W}}^{2}\}\right)\leq\delta$ (A.2)
@@ A.1.1 Chaining Argument
[…]
@@
First, by Lemma 3, with an inner probability at least $21-\delta/$ 2:
@@ A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma 3)
$\begin{split}U(\tau)&\leq\mathbb{E}^{*}\left[\exp\left(2\tau\cdot\sup_{z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}}~{}\max_{v\in V_{1/2}}\langle W_{m}(z_{1},t),v\rangle\right)\right]\\ \end{split}$ (A.10)
@@
In this product probability space, the sample $\xi_{1:m}$ corresponds to the first $m$ coordinates projections and the additional Rademacher random variables $\varepsilon_{1:m}$ depending only on the $(m+1)$st coordinate. The outer integral $\mathbb{E}^{*}$ in equation (A.11) then refers to the outer integral for this probability space.
@@ Appendix B Proof of Proposition 8
There exist a numerical constant $C>0$ so that for $m\geq Cd$, with an inner probability at least $1-2\exp(-d)$ [^1] The original expression in the paper [14], $1-2\exp(-\min\{c_{2}m,d^{2}\})$, should be corrected to $1-2\exp(-d)$. , the set of stationary points $\mathcal{Z}_{S}$ of $\Phi_{S}$ satisfies:
@@ Appendix C Proof of Corollary 2
For the nonsmooth function $h(z)=|z|$, the corresponding value of $\zeta$ is given by $\zeta=3$ as 00 is the only nondifferentiable point of $h$. The corresponding $\mathcal{F}$ is given by
=$\mathcal{F}=\left\{\{u\in\mathbb{R}^{d_{1}},v\in\mathbb{R}^{d_{2}},b\in\mathbb{R}:\langle u,y\rangle\langle v,w\rangle-b\geq t\}\mid y\in\mathbb{R}^{d_{1}},w\in\mathbb{R}^{d_{2}},t\in\mathbb{R}\right\}.$
@@ Appendix D Proof of Corollary 3
For the nonsmooth function $h(z)=|z|$, the corresponding value of $\zeta$ is given by $\zeta=3$ as 00 is the only nondifferentiable point of $h$. The corresponding $\mathcal{F}$ is given by
=$\mathcal{F}=\left\{\{A\in\mathbb{R}^{D\times D},b\in\mathbb{R}:\langle A,XX^{T}\rangle-b\geq t\}\mid X\in\mathbb{R}^{D\times r_{0}},t\in\mathbb{R}\right\}.$

### v6

@@ 1 Introduction
We instantiate this principle in the context of stochastic weakly convex objectives, a broad class of nonsmooth, nonconvex functions encompassing many widely studied models [3, 4, 5]. Our main application is to stochastic convex-composite objectives [6, 7], which include robust phase retrieval [8], blind deconvolution [9], matrix sensing [10, 11], conditional Value-at-Risk [12], and nonlinear quantile regressions [13]. These models appear across imaging, statistics, machine learning, and risk management. For this class, we derive uniform convergence rates for subdifferentials by studying the convergence behavior of stochastic subgradient mappings (Theorem 4 and Theorem 5). Our analysis complements prior work by (i) removing differentiability assumptions on the population risk, and (ii) establishing tight convergence rates under the Hausdorff distance, a topologically weaker notion than graphical convergence that more directly captures the geometry of subdifferential sets. For instance, in the setting of robust phase retrieval [8, 14], we establish a sharp convergence rate of $\sqrt{d/m}$ (modulo logarithmic factors on $d,m$) for the uniform convergence of subdifferential mappings under the Hausdorff distance, where $d$ is the problem’s dimension and $m$ is the sample size (Corollary 1). This is complementary to the state-of-the-art $\sqrt[4]{d/m}$ rate, which is established under the graphical distance of subdifferentials using an alternative approach based on Attouch’s epigraphical convergence theorem [15, 16]. Our results lead to new understanding into nonsmooth empirical risk landscapes, including how the locations of stationary points evolve with sample size—providing a sharper understanding of finite-sample behavior in high-dimensional stochastic optimization.
@@ 2 Notation and Basic Definitions
For $x,y\in\mathbb{R}$, we let $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$. In $\mathbb{R}^{d}$, we use $\left\|{\cdot}\right\|$ and $\langle\cdot,\cdot\rangle$ to denote the standard Euclidean norm and inner product respectively. We let $\mathbb{B}(x_{0};r)=\{x:\left\|{x-x_{0}}\right\|<r\}$ denote the open $\ell_{2}$ ball with center at $x_{0}\in\mathbb{R}^{d}$ and radius $r$. For matrices in $\mathbb{R}^{d\times d}$, we identify each such matrix with a vector in $\mathbb{R}^{d^{2}}$ and define $\langle A,B\rangle$ as ${\rm tr}(AB^{T})$, where $\rm tr$ denotes the trace. For a set $X$, we use ${\rm cl},{\rm int}$ to denote the closure and interior respectively for the set $X$. For a closed convex set $X$, we use $\iota_{X}$ to denote the $+\infty$-valued indicator for the set $X$, that is $\iota_{X}(x)=0$ if $x\in X$ and $\infty$ otherwise. The normal cone to $X$ at $x$ is $\mathcal{N}_{X}(x):=\{v\in\mathbb{R}^{d}:\langle v,y-x\rangle\leq 0~\text{for all}~y\in X\}$.
@@ 2.2 Set-valued Analysis
For a set $A\subseteq\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$, we denote by ${\rm dist}(y,A)=\inf_{z\in A}\left\|{y-z}\right\|$ the distance from $y$ to $A$ with respect to the Euclidean norm $\left\|{\cdot}\right\|$. For two sets $A_{1},A_{2}\subseteq\mathbb{R}^{d}$, we denote by
@@ 2.4 Measurability Issues
[…]
@@ 3.1 Uniform Subdifferential Bounds via Selections
[…]
@@
We now argue that we can, without loss of generality, assume that for every $k$, $x_{k}\in D_{2}$ (the set of differentiable points of $f_{2}$) as well. If some $x_{k}\not\in D_{2}$, applying Clarke’s subdifferential characterization to $f_{1}$ at $x_{k}$ (3.2), we can replace such $x_{k}$ with some $x_{k}^{\prime}$ belonging to the subset $D_{1}\cap D_{2}$ dense in $\mathcal{O}$ (note $\mathcal{O}\backslash(D_{1}\cap D_{2})$ has Lebesgue measure 0) such that $\left\|{x_{k}^{\prime}-x_{k}}\right\|\to 0$ and $\left\|{\nabla f_{1}(x_{k}^{\prime})-\nabla f_{1}(x_{k})}\right\|\to 0$ as $k\to\infty$. As a result, $x_{k}^{\prime}$ preserves all the properties of $x_{k}$, including $x_{k}^{\prime}\to x$, $x_{k}^{\prime}\in D_{1}$ and $\nabla f_{1}(x_{k}^{\prime})\to v_{1}^{*}$, while guaranteeing that $x_{k}^{\prime}\in D_{2}$ for all $k$.
@@ 3.3 A Class of Stochastic Weakly Convex Minimization Problems
We now revisit the conditions on the convex function $h$ and smooth function $c(x,\xi)$ that guarantees the weak-convexity assumptions for the composition $f(x,\xi)=h(c(x,\xi))$, ensuring the regularity required for subdifferential calculus rules. We take the following conditions—local Lipschitzian and integrability conditions on h,ch,c—from [7, Section 2.2].
@@ 4.1 Subgradient Selections
where $z\mapsto\mathbf{1}(z\geq t_{j})$ is an indicator function that is a subgradient for the mapping $z\mapsto(z-t_{j})_{+}$, taking the value $1$ when $z\geq t_{j}$ and 0 otherwise. Combining these, we define a subgradient $g$ for $h$ at every $z$:
@@ 4.2 A Uniform Bound on Subgradient Selections
$\mathcal{F}=\{\{\xi\in\Xi\colon c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}.$ (4.12)
@@ 4.3 Implication: A Uniform Bound on Subdifferentials
In the literature on uniform convergence rates of subdifferential mappings for stochastic weakly-convex objectives, methods using Moreau envelope smoothing [26] and Attouch’s epigraphical convergence theorem [15] achieve a convergence rate of $\sqrt[4]{1/m}$ suboptimal in terms of the sample size $m$ [14, 16]. This $\sqrt[4]{1/m}$ convergence is measured using graphical distance, a notion which is topologically stronger than the uniform convergence criterion we consider in Theorem 5 (see Remark of Theorem 1). Our work complements these studies by showing a convergence rate’s dependence on sample size $m$ to be $\sqrt{1/m}$ (modulo logarithmic factors) under a topologically weaker notion of convergence.
@@ 5 VC Dimension Bounds
$\mathcal{F}=\{\{\xi\in\Xi\colon c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}$
@@
$\begin{split}\mathcal{B}&=\{(b_{1},b_{2},\ldots,b_{N})\in\{0,1\}^{N}\colon\\ &~~~~~~~~~~\text{there exists $z\in\mathbb{R}^{d+1}$ such that}~b_{i}=\mathbf{1}\{q_{i}(z)\geq 0\}~\text{for all $1\leq i\leq N$}\}.\end{split}$
@@ 6.1 Illustration I: Phase Retrieval
as 0 is the only nondifferentiable point of $h$. The corresponding $\mathcal{F}$ is given by
$\mathcal{F}=\left\{\{a\in\mathbb{R}^{d},b\in\mathbb{R}\colon\langle a,x\rangle^{2}-b\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\right\}.$
@@ 6.1.1 Noiseless Phase Retrieval
For standard normal measurement vectors $a\sim\mathsf{N}(0,I)$, a recent important and interesting study first characterizes $\mathcal{Z}$, the locations of stationary points of the population objective $\Phi(x)$ [14]. Remarkably, the authors further develop a quantitative version of Attouch’s epi-convergence theorem to show that the stationary points of the finite samples objective $\Phi_{S}$ converge to those of the population objective $\Phi$ at a rate of $\sqrt[4]{d/m}$ [14]. More precisely, [14, Theorem 5.2, Corollary 6.3] prove that, there exist numerical constants $c,C>0$ such that when $m\geq Cd$, with an inner probability at least $1-C\exp(-cd)$ (recall the definition of inner probability from Section 2.4):
@@ 7 Discussion
  • Can Theorem 1 be extended to other function classes? An intriguing candidate is the class of difference-of-convex functions [58].
@@
Acknowledgements Feng Ruan extends his gratitude to his friends and advisor—John Duchi, Qiyang Han, X.Y. Han, Kate Yao and Yichen Zhang—for their support and patience as he tackled this once seemingly challenging problem. A shift in perspective during his time at Northwestern led him to this solution. He is also grateful to Damek Davis and Dmitriy Drusvyatskiy for their valuable feedback on the initial manuscript draft posted on arXiv. Feng Ruan would also like to express gratitude to Johannes Royset, Alexander Shapiro and Lai Tian, and the anonymous reviewers for their valuable discussions on literature and insights regarding the proof of Theorem 1 using variational analytic tools and also its limitations.
@@ Appendix A Proof of Theorem 4
$\mathbb{P}^{*}\left(\sup_{z_{1}\in\mathcal{Z}_{1},z_{2}\in\mathcal{Z}_{2},t\in\mathbb{R}}\left\|{K_{m}(z_{1},z_{2},t)-K(z_{1},z_{2},t)}\right\|\geq C(\sigma r+\sigma_{0})\cdot\max\{\Delta_{\mathcal{W}},\Delta_{\mathcal{W}}^{2}\}\right)\leq\delta$ (A.2)
@@ A.1.1 Chaining Argument
This decomposition is valid because the limit $R_{m}(z_{1},z_{2},t)=\lim_{k\to\infty}R_{m}(z_{1},\pi_{k}(z_{2}),t)$ holds with probability one, as $z_{2}\mapsto R_{m}(z_{1},z_{2},t)$ is continuous by assumption. Applying the supremum over $z_{1}$, $z_{2}$, $t$ after taking the norm $\left\|{\cdot}\right\|$, and using the triangle inequality we obtain:
@@
First, by Lemma 3, with an inner probability at least $1-\delta/2$:
@@ A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma 3)
[…]
@@
In this product probability space, the sample $\xi_{1:m}$ corresponds to the first $m$ coordinates projections and the additional Rademacher random variables $\varepsilon_{1:m}$ depending only on the (m+1)(m+1)st coordinate. The outer integral $\mathbb{E}^{*}$ in equation (A.11) then refers to the outer integral for this probability space.
@@ Appendix B Proof of Proposition 8
There exist a numerical constant $C>0$ so that for $m\geq Cd$, with an inner probability at least $2\exp(-d)11$1The original expression in the paper [14], $1-2\exp(-\min\{c_{2}m,d^{2}\})$, should be corrected to $1-2\exp(-d)$. , the set of stationary points $\mathcal{Z}_{S}$ of $\Phi_{S}$ satisfies:
@@ Appendix C Proof of Corollary 2
For the nonsmooth function $h(z)=|z|$, the corresponding value of $\zeta$ is given by $\zeta=3$ as 0 is the only nondifferentiable point of $h$. The corresponding $\mathcal{F}$ is given by
$\mathcal{F}=\left\{\{u\in\mathbb{R}^{d_{1}},v\in\mathbb{R}^{d_{2}},b\in\mathbb{R}:\langle u,y\rangle\langle v,w\rangle-b\geq t\}\mid y\in\mathbb{R}^{d_{1}},w\in\mathbb{R}^{d_{2}},t\in\mathbb{R}\right\}.$
@@ Appendix D Proof of Corollary 3
For the nonsmooth function $h(z)=|z|$, the corresponding value of $\zeta$ is given by $\zeta=3$ as 0 is the only nondifferentiable point of $h$. The corresponding $\mathcal{F}$ is given by
$\mathcal{F}=\left\{\{A\in\mathbb{R}^{D\times D},b\in\mathbb{R}:\langle A,XX^{T}\rangle-b\geq t\}\mid X\in\mathbb{R}^{D\times r_{0}},t\in\mathbb{R}\right\}.$